### v2

# Garf: Learning Generalizable 3D Reassembly for Real-World Fractures

Abstract
3D reassembly is a challenging spatial intelligence task with broad applications across scientific domains. While large-scale synthetic datasets have fueled promising learning-based approaches, their generalizability to different domains is limited. Critically, it remains uncertain whether models trained on synthetic datasets can generalize to real-world fractures where breakage patterns are more complex. To bridge this gap, we propose Garf, a generalizable 3D reassembly framework for real-world fractures. Garf leverages fracture-aware pretraining to learn fracture features from individual fragments, with flow matching enabling precise 6-DoF alignments. At inference time, we introduce two-session flow matching, improving robustness to unseen objects and varying numbers of fractures. In collaboration with archaeologists, paleoanthropologists, and ornithologists, we curate Fractura, a diverse dataset for vision and learning communities, featuring real-world fracture types across ceramics, bones, eggshells, and lithics. Comprehensive experiments have shown our approach consistently outperforms state-of-the-art methods on both synthetic and real-world datasets, achieving 82.87% lower rotation error and 25.15% higher part accuracy on the Breaking Bad Everyday dataset. This sheds light on training on synthetic data to advance real-world 3D puzzle solving, demonstrating its strong generalization across unseen object shapes and diverse fracture types. Garf’s code, data and demo are available at https://ai4ce.github.io/GARF/.
Figure 1: We curate Fractura, a unique dataset presenting real-world fracture assembly challenges across scientific domains, including ceramics, bones, eggshells, and lithics. To tackle these challenges, we introduce Garf, a generalizable 3D reassembly framework designed to handle varying object shapes, diverse fracture types, and the presence of missing or extraneous fragments.
footnotetext: ✉Corresponding authors {z.jing, cfeng}@nyu.edu.
1 Introduction
We have long been captivated by questions of human origins [wilson2012social]: What are we? Where do we come from? The answers to these fundamental questions lie in archaeological materials such as bones [callaway2017oldest], ceramics [wang2019emergence], and lithics [mcpherron2010evidence]. However, these artifacts are often highly fragmented and incomplete [callaway2017oldest, laughlin_149_2005]. Reassembling them requires placing each fragment in its correct position and orientation to restore a complete or functional entity [scarpellini2024diffassemble]. This process is not only time-consuming but also challenges the limits of human spatial intelligence [scarpellini2024diffassemble, gardner2011frames, tsesmelis2025re]. For instance, an experimental study on lithic refitting reported a success rate of only 30%, with experts performing only marginally better than novices [laughlin_experimental_2010]. Consequently, museum storerooms around the world remain filled with thousands of unassembled fragments, waiting to be pieced back together [tsesmelis2025re].
Recently, the emergence of the large-scale dataset Breaking Bad [sellan2022breaking] has brought new hope to this domain, fueling the development of data-intensive reassembly methods [wu2023leveraging, villegas2023matchmakernet, zhang2024scalable, li2024geometric, lu2024jigsaw, cui2024phformer, scarpellini2024diffassemble, wang2024puzzlefusion++, lu2024jigsaw++, xu2024fragmentdiff, lee20243d, kim2024fracture]. While PuzzleFusion++ achieves state-of-the-art (SOTA) performance on the everyday subset, its generalization degrades significantly on the artifact subset due to its reliance on global geometry learning [wang2024puzzlefusion++]. More critically, the Breaking Bad dataset is generated via physics-based simulation [sellan2023breaking], raising a fundamental question: can models trained on synthetic breakage patterns generalize to real-world fractures?
Figure 2: Characteristics and Challenges of Fractura. (i) Diverse fracture types including two synthetic and three real-world types, across ceramics, bones, eggshells, and lithics. (ii) Real-world challenges such as missing or extraneous fragments.
To answer this question, we identify two major real-world fracture challenges and curate Fractura, a dataset capturing key complexities: (i) Data diversity encompasses three geometrically distinct fracture types across multiple scientific domains, including bones, ceramics, eggshells, and lithics. As shown in Fig. 2, ceramics exhibit irregular, chaotic fractures typical of random breakage events, whereas flintknapping produces conchoidal fractures. This allows a systematic study of how object shapes and fracture types affect reassembly performance. (ii) Missing or extraneous fragments are common issues in real-world scenarios [harvati2019apidima, tsesmelis2025re] (see Fig. 2), which guides our model design to improve robustness.
In response to these complexities, we propose Garf, a generalizable 3D reassembly framework for real-world fractures, featuring four components: (i) Large-scale fracture-aware pretraining takes lessons from recent successes of large-scale pretraining in natural language and computer vision [oquab2023dinov2, radford2021learning, stevens2024bioclip, achiam2023gpt, steiner2024paligemma]. This module learns fracture segmentation of individual fragments to handle unseen objects as well as missing or extraneous fragments. (ii) Flow-based reassembly on $\mathrm{SE}(3)$ introduces flow matching to learn pose distribution, leveraging the $\mathrm{SO}(3)$ manifold for accurate rotation estimation. Inspired by human puzzle-solving, we design a multi-anchor training strategy that randomly selects a subset of fragments to form local structures, exposing the model to diverse combinations to enhance distribution learning. (iii) Two-session flow matching at inference time emphasizes the importance of pose initialization. The first session performs a one-step flow matching to generate a strong pose initialization on SE(3); the second session performs standard multi-step flow for refinement. (iv) LoRA-based fine-tuning enhances the model’s adaptability to specific domains.
Our main contributions are as follows:
  • We introduce Garf, the first flow-based 3D fracture assembly framework that integrates fracture-aware pretraining and two-session flow matching, achieving SOTA performance across synthetic and real-world datasets.
  • Garf sheds light on training on synthetic data for real-world challenges, effectively handling unseen object shapes and diverse fracture types, while remaining robust to missing or extraneous fragments.
  • Collaborating with domain experts, we curate Fractura, a diverse dataset capturing real-world fracture complexities, to conduct the first study on 3D reassembly generalization across ceramics, bones, eggshells, and lithics. Moreover, we apply the first integration of LoRA-based fine-tuning for domain-adaptive 3D reassembly.
2 Fractura Dataset
Existing datasets, whether synthetic [sellan2022breaking] or real [lamb2023fantastic, tsesmelis2025re], are limited to a single fracture type. To fill this gap, we curate Fractura, a challenging dataset (see Fig. 2), designed for a comprehensive evaluation of how object shapes, fracture types, and the presence of missing or extraneous parts affect reassembly performance.
Data Characteristics. Fractura comprises both real and synthetic fracture subsets. In collaboration with archaeologists, paleoanthropologists, and ornithologists, the real fracture subset includes three real fracture types relevant to scientific challenges (see Fig. 2): (i) Random breakage produces irregular, chaotic fractures, commonly observed in ceramics, bones, and eggshells. (ii) Incomplete ossification results in unfused bone ends (epiphyses) in juvenile skeletons, leading to fragmented rather than intact bones in skeletal collections and fossil records. Reassembling these unfused parts will benefit the following analysis and further create a complete series of bone developments over time from early childhood to adulthood. (iii) Flintnapping produces conchoidal fractures in lithics, characterized by radially propagating fracture lines (see Fig. 2). Real-world collections and scans naturally incorporate missing or extraneous fragments. For the synthetic fracture subset, we generate realistic fractures using physics-based simulation [sellan2023breaking] for ceramics, bones, and eggshells, and geometry-based simulation [orellana2021proof] for lithics.
Data Collection and Simulation. We utilize the high-accuracy Artec spider 3D scanner (precision: 0.05 mm) to acquire detailed 3D meshes of real fragments and intact objects from the same categories. The real fracture subset serves as test data, while intact objects are used to generate the synthetic fracture subset for fine-tuning and evaluation. Details are provided in the supplementary materials.
Table 1: Comparisons of 3D Reassembly Datasets. For Fracture Type*: Syn. denotes synthetic data. The number in parentheses indicates the number of synthetic/real fracture modes.
Datasets Breaking Fantastic RePAIR [tsesmelis2025re] Fractura
Bad [sellan2022breaking] Breaks [lamb2023fantastic]
# Pieces 8M 300 1070 53350+292
# Breaks 2-100 2 2-44 2-22
# Assemblies 10474 150 117 9727+41
Fracture Type* Syn. (1) Real (1) Real (1) Syn. (2) + Real (3)
Object Type Everyday Everyday Frescoes Bones, Eggshells
Artifact, Other Lithics, Ceramics
Miss. / Extra. $\times$ / $\times$ $\times$ / $\times$ $\checkmark$ / $\times$ $\checkmark$ / $\checkmark$
Texture $\times$ $\times$ $\checkmark$ $\checkmark$
Data Statistics. Table 1 summarizes key statistics of Fractura and other 3D reassembly datasets. Compared with existing datasets, Fractura introduces a diverse set of real and synthetic fractures across multiple scientific domains. We are actively expanding its size and diversity. More details are provided in the supplementary materials.
3 Method
Previous work either enhances global geometry learning through fragment features [wu2023leveraging, scarpellini2024diffassemble, wang2024puzzlefusion++] or jointly learns hierarchical features from both global and local geometry [lu2024jigsaw]. In contrast, as shown in Fig. 3, Garf decouples the understanding of local fracture features (Sec. 3.1) and global fragment alignment (Sec. 3.2). At inference, we propose the two-session flow matching (Sec. 3.3) for robustness to unseen objects and increasing numbers of fractures. To further boost the performance on domain-specific data, we employ a LoRA-based fine-tuning method (Sec. 3.4).
3.1 Why Large-Scale Fracture-Aware Pretraining?
Figure 3: Pipeline of Garf. Our framework comprises two main components: (i) Fracture-aware pretraining leverages $14\times$ more data than previous methods to learn the local fracture features via fracture point segmentation, and (ii) Two-session flow-based reassembly on $\mathrm{SE}(3)$ leverages the $\mathrm{SO}(3)$ manifold for precise rotation estimation. At inference time, one-step pre-assembly strategy provides better initial poses, enhancing robustness against unseen objects and increasing numbers of fractures.
Humans can infer fracture points without knowledge of the global object shape [papaioannou2001virtual, lu2024jigsaw]. To emulate this ability, we leverage large-scale data to learn local fracture features from individual fragments, drawing inspiration from recent advances in large-scale pretraining [oquab2023dinov2, radford2021learning, stevens2024bioclip, achiam2023gpt, steiner2024paligemma]. This enables our method to generalize to unseen object shapes and handle missing or extraneous fragments.
Specifically, we sample a set of point clouds $\bm{P}=\{P^{1},P^{2},\dots,P^{N}\}$ to represent fragments, where $N$ is the number of fragments. To extract fracture-level features from $\bm{P}$, we employ Point Transformer V3 (PTv3) [wu2024point] as the backbone, integrating two MLP layers [lu2024jigsaw] as the segmentation head for fracture cloud segmentation. Since fracture points constitute only a small proportion of $\bm{P}$, an imbalance arises between positive and negative samples. To mitigate this, we adopt the Dice loss function [sudre2017generalised]:
$\mathcal{L}_{\text{Seg}}=1-\frac{2\sum_{i=1}^{N}p_{i}g_{i}+\epsilon}{\sum_{i=1}^{N}p_{i}+\sum_{i=1}^{N}g_{i}+\epsilon},$ (1)
where $p$ is the predicted value and $g$ is its ground truth label. To derive the high-quality $g$, we directly extract shared surfaces between connected fragments from the mesh, defining them as fracture surfaces $\bm{F}$. We then apply Poisson disk sampling [bridson2007fast] to generate $\bm{P}$, ensuring that $g=\bm{P}\cap\bm{F}$. This weighted sampling based on the surface area of fragments prevents the encoder from overfitting to specific point densities, improving generalization.
3.2 Flow-Based Reassembly on SE(3)
Inspired by flow-based generative models [lipman2022flow, liu2022flow, pooladian2023multisample, lipman2024flow] in image synthesis [esser2024scaling], protein structure generation [bose2023se, huguet2024sequence, geffnerproteina], and robotic manipulation [black2024pi0, yim2023fast], we leverage flow matching (FM) to model the fracture reassembly process.
On a manifold $\mathcal{M}$, the flow $\psi_{t}:\mathcal{M}\rightarrow\mathcal{M}$ is defined as the solution of an ordinary differential equation (ODE):
$\frac{\mathrm{d}}{\mathrm{d}t}\psi_{t}(\bm{x})=\bm{v}_{t}(\psi_{t}(\bm{x})),\quad\psi_{0}(\bm{x})=\bm{x},$ (2)
where $\bm{v}_{t}(\bm{x})\in\mathcal{T}_{\bm{x}}\mathcal{M}$ is the time-dependent vector field, and $\mathcal{T}_{\bm{x}}\mathcal{M}$ denotes the tangent space at $\bm{x}$. In the context of $\mathrm{SE}(3)$, the tangent space corresponds to the Lie algebra $\mathfrak{se}(3)$, a six-dimensional vector space representing the velocity of the rigid motion of fragments.
Given an object composed of $N$ fragments, we represent their poses as $\bm{T}=\{T^{1},T^{2},\cdots,T^{N}\}$, where each $T^{i}$ consists of a rotation $r\in\mathrm{SO}(3)$ and a translation $a\in\mathbb{R}^{3}$, expressed as $T^{i}:\{r,a\}\in\mathrm{SE}(3)$. The initial noise distribution is defined as: $p_{0}(\bm{T}_{0})=\mathcal{U}(\mathrm{SO}(3))\otimes\mathcal{N}(0,I_{3})$, where the rotation noise follows a uniform distribution over $\mathrm{SO}(3)$, and the translation noise is sampled from a unit Gaussian distribution. We decouple the rotation and translation flows, allowing independent flow modeling in $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$. Therefore, the conditional flow $\bm{T}_{t}=\psi_{t}(\bm{T}_{0}\mid\bm{T}_{1})$ follows the geodesic path connecting $\bm{T}_{0}$ and $\bm{T}_{1}$:
$\bm{r}_{t}$ $=\exp_{\bm{r}_{0}}(t\log_{\bm{r}_{0}}(\bm{r}_{1})),$ (3)
$\bm{a}_{t}$ $=(1-t)\bm{a}_{0}+t\bm{a}_{1},$
where $\exp_{\bm{r}}$ and $\log_{\bm{r}}$ are the exponential and logarithmic maps on $\mathrm{SO}(3)$. The final optimization objective is:
$\mathcal{L}_{\text{FM}}=$ $\mathbb{E}_{t,p_{1}(\bm{T}_{1}),p_{t}(\bm{T}|\bm{T}_{1})}\Biggl[\sum_{i=1}^{N}\Bigl\|\bm{v}^{i}_{r}(\bm{T}_{t},t)-\frac{\log_{\bm{r}_{t}}(\bm{r}_{1})}{1-t}\Bigr\|^{2}_{g}$
$+\Bigl\|\bm{v}^{i}_{a}(\bm{T}_{t},t)-\frac{\bm{a}_{1}-\bm{a}_{t}}{1-t}\Bigr\|^{2}_{g}\Biggr].$ (4)
Further details are in the supplementary materials.
Network Architecture. Consider an object consisting of $N$ fragments with inital poses $\bm{T}_{t}\in\mathbb{R}^{N\times 7}$ at timesteps $\bm{t}$. The corresponding latent features are extracted from the pre-trained encoder: $\bm{F}=\mathcal{E}(\bm{P})\in\mathbb{R}^{M\times c}$, where $c$ is the number of channels and $M$ is the pre-defined number of sampled points. We integrate point cloud coordinates $\bm{P}\in\mathbb{R}^{M\times 3}$, normals $\bm{n}\in\mathbb{R}^{M\times 3}$, and scale information $\bm{s}\in\mathbb{R}^{M\times 1}$ as pose-invariant shape priors in the position embedding:
$\bm{s}_{\text{emb}}=f_{\text{shape}}\Big(\operatorname{concat}\Big(\bm{F},\,\texttt{PE}(\bm{P}),\,\texttt{PE}(\bm{n}),\,\texttt{PE}(\bm{s})\Big)\Big).$ (5)
The pose is treated as spatial information: $\bm{p}_{\text{emb}}=f_{\text{pose}}\big(\texttt{PE}(\bm{T})\big)$. These embeddings are combined to form the Transformer input: $\bm{d}=\texttt{PE}(\bm{s}_{\text{emb}}+\bm{p}_{\text{emb}})$.
The feature $\bm{d}$ is then processed through $L$ Transformer layers including self-attention and global attention. Leveraging the efficient computation of FlashAttention [dao2023flashattention2], we modify the self-attention module to allow more attention on large fragments:
$\begin{gathered}\bm{Q},\,\bm{K},\,\bm{V}=\operatorname{Linear}(\operatorname{LN}(\bm{d})),\\ \bm{A}_{\text{self}}=\operatorname{FlashAttn}\Big(\bm{Q},\,\bm{K},\,\bm{V};\,\operatorname{cu}(\ell),\,\ell_{\max}\Big),\end{gathered}$ (6)
where the model computes the sequence lengths $\ell$ of the variable number of sampled points per fragment batch, along with their cumulative sums $\operatorname{cu}(\ell)$. More details are in the supplementary materials. The self-attention output $\bm{A}_{\text{self}}$ is added to the original feature $\bm{d}$, yielding the updated representation $\bm{h}=\bm{d}+\bm{A}_{\text{self}}$. The global attention layer then applies a similar attention mechanism to aggregate information across fragments with $\ell=M$. The output undergoes Layer Normalization (LN) and a feed-forward network (FFN) with a residual connection, before regressing the pose $\bm{T}_{t-1}:\{\bm{r}_{t-1},\bm{a}_{t-1}\}$ at the next timestep:
$\begin{gathered}\bm{h}\leftarrow\bm{h}+\operatorname{FFN}(\operatorname{LN}(\bm{h})),\\ \bm{a}_{t-1}=f_{\text{trans}}(\bm{h}),\quad\bm{r}_{t-1}=f_{\text{rot}}(\bm{h}).\end{gathered}$ (7)
Multi-Anchor Training Strategy. We observe that probability paths vary significantly across fragments; some require complex transformations, while others remain nearly stationary. To model this variation, we introduce a multi-anchor training strategy, randomly selecting $k\in[1,N-1]$ fragments as anchors and fixing their positions with identity rotations and zero translations. For these anchor fragments $\bm{i}$, the vector field is explicitly supervised to be zero: $\bm{v}^{\bm{i}}(\bm{T}_{t},t)=0$. Unlike prior approaches [wang2024puzzlefusion++], which prevent gradient propagation for a single anchor (the largest fragment and its connected neighbors within a 50% threshold), our multi-anchor strategy enforces a zero vector field, expanding the range of probability paths and enhancing generalization across diverse fragment configurations.
3.3 Two-Session Flow Matching at Inference Time
Prior diffusion/flow matching work only uses single-session inference. While image generation methods use search strategies or self-supervised verifiers to improve initialization [ma2025inference, zhou2024golden, li2024enhancing, qi2024not], such strategies are less effective for pose estimation. In contrast, we design the two-session flow matching at inference time: the first session performs a one-step FM inference ($\text{step}=1$) to generate an initial pose $\bm{T}_{0}^{\prime}$, leveraging FM’s ability to model straight-line probability paths and narrow the search space; the second session performs standard multi-step flow for refinement. Despite a minimal 5% increase in computational cost ($20\rightarrow 1+20$ steps), this design outperforms single-session inference, particularly for larger fragment sets.
Figure 4: Qualitative Comparisons on the Breaking Bad and Fractura. Garf consistently produces more accurate reassemblies, particularly on the Breaking Bad Artifact subset and Fractura synthetic fracture subset, demonstrating strong generalization to unseen object shapes. Meshes are used for visualization only. Additional results are available in the supplementary material.
3.4 LoRA-based Fine-tuning
To quickly adapt to domain-specific contexts, we employ a LoRA-based [hu2022lora] fine-tuning approach using the synthetic fracture subset in Fractura. Specifically, we integrate LoRA adapters into the self-attention and global attention layers of the final Transformer block while unfreezing the MLP heads for pose prediction ($f_{\text{rot}}$ and $f_{\text{trans}}$). Our experiments demonstrate that this lightweight fine-tuning method requires as few as 5–10 domain-specific objects to achieve substantial improvements in scientific applications such as juvenile skeleton reconstruction and lithic refitting.
4 Experiments
4.1 Training and Evaluation Details.
[…]
Datasets. We evaluate our model on three datasets with diverse object shapes and fracture types: (i) Breaking Bad [sellan2022breaking], the largest synthetic fracture dataset for 3D reassembly. We use the volume-constrained version, evaluating 7,872 assemblies from the Everyday subset and 3,697 assemblies from the Artifact subset. Results on the vanilla version are in the supplementary materials. (ii) Fantastic Breaks [lamb2023fantastic], a real-world dataset of 195 manually scanned fractured objects with complex surfaces, used for evaluation only. (iii) Fractura, a mixed synthetic-real dataset spanning ceramics, bones (vertebrae, limbs, ribs), eggshells, and lithics. For the synthetic fracture subset, we follow an 80/20 split [tsesmelis2025re] for LoRA fine-tuning and evaluation.
Evaluation Metrics. Following [wang2024puzzlefusion++], we evaluate assembly quality using four metrics: (i) RMSE(R) is the root mean square error of rotation (degrees); (ii) RMSE(T) is the root mean square error of translation; (iii) PA is the percentage of correctly assembled fragments, where the per-fragment chamfer distance is below 0.01; and (iv) CD is the chamfer distance between the assembled object and ground truth.
Competing Methods. We compare our approach against Global [li2020learning], LSTM [wu2020pq], DGL [zhan2020generative], Jigsaw [lu2024jigsaw], PMTR***The performance of PMTR [lee20243d] is from PuzzleFusion++ [wang2024puzzlefusion++]. [lee20243d], and PuzzleFusion++ (PF++)  [wang2024puzzlefusion++]. We implemented PF++ [wang2024puzzlefusion++] and Jigsaw [lu2024jigsaw] using its official codebase, while performance metrics for other methods on the volume-constrained Breaking Bad dataset are sourced from their official papers or repositories. Additional comparisons with SE(3)-Equiv [wu2023leveraging], DiffAssemble [scarpellini2024diffassemble], and PHFormer [cui2024phformer] on the vanilla Breaking Bad dataset, as well as FragmentDiff [xu2024fragmentdiff] on its custom Breaking Bad subset, are provided in the supplementary materials.
4.2 Can Garf Generalize to Unseen Shapes?
Real-world fragmentation varies in object geometries, requiring Garf to generalize to unseen object shapes. To empirically assess this, we evaluate Garf-mini on Fractura and the Artifact subset of Breaking Bad, as well as Garf on Fractura.
Results and Analysis. Table 2 presents the evaluation results on the Breaking Bad dataset, where Garf achieves significant improvements over previous SOTA methods. Compared to PF++ [wang2024puzzlefusion++], Garf reduces rotation error by 82.87% and translation error by 79.83%, while achieving a CD below 0.001, indicating near-perfect reconstructions. Even more impressively, Garf-mini demonstrates exceptional generalization capability. Despite being trained only on the Everyday subset, it maintains consistent performance on the unseen Artifact subset, avoiding the performance degradation typically observed with unseen object shapes. This highlights Garf’s robust feature extraction and assembly mechanisms. Notably, Garf achieves 95.33% and 95.04% PA on the Breaking Bad dataset, approaching the theoretical maximum PA of 96.49% (Everyday subset) and 96.10% (Artifact subset) when excluding fragments smaller than 0.1% of the object volumeScanning tiny 3D fragments in real-world settings is inherently challenging; anthropologists often treat such fragments as missing parts.. On the challenging synthetic subset of Fractura, which contains unseen domain-specific objects, Garf further demonstrates superior generalization capability, outperforming all competing methods across all evaluation metrics. As shown in Fig. 4, Garf consistently produces more accurate reassemblies, particularly on the Breaking Bad Artifact subset and the Fractura synthetic fracture subset, confirming its strong generalization to unseen object shapes.
Table 2: Quantitative Results on Volume-Constrained Breaking Bad [sellan2022breaking] and Fractura Datasets. The best performance metric is highlighted in bold, while the second-best is underlined.
Methods RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$ CD $\downarrow$
degree $\times 10^{-2}$ % $\times 10^{-3}$
Tested on the Everyday Subset
Global [li2020learning] 80.50 14.60 28.70 13.00
LSTM [wu2020pq] 82.70 15.10 27.50 13.30
DGL [zhan2020generative] 80.30 13.90 31.60 11.80
Jigsaw [lu2024jigsaw] 42.19 6.85 68.89 8.22
PMTR [lee20243d] 31.57 9.95 70.60 5.56
PF++ [wang2024puzzlefusion++] 35.61 6.05 76.17 2.78
Garf-mini 6.68 1.34 94.77 0.25
Garf 6.10 1.22 95.33 0.22
Tested on the Artifact Subset
Jigsaw 43.75 7.91 65.12 8.50
PF++ 47.03 10.63 57.97 8.24
Garf-mini 7.67 1.77 93.34 0.81
Garf 5.82 1.27 95.04 0.42
Tested on the Fractura (Synthetic Fracture)
Jigsaw 60.50 18.49 33.06 70.68
PF++ 62.57 18.65 37.74 36.13
Garf-mini 27.88 6.79 76.25 7.70
Garf 19.63 4.93 83.41 6.06
4.3 How Do Fracture Types Affect Generalization?
As scanning time significantly increases with fragment count [lamb2023fantastic], real fracture datasets are typically limited in size and diversity, making large-scale training infeasible. To address this, we investigate how fracture types impact zero-shot generalization on Fantastic Breaks and Fractura, as well as the fine-tuning performance using domain-specific synthetic fractures on Fractura.
Real Fracture on Fantastic Breaks [lamb2023fantastic]. As shown in Table 3, Garf demonstrates superior generalization to real-world fracture surfaces, achieving a remarkable 48.65% reduction in rotation error compared to PF++ [wang2024puzzlefusion++], indicating that our model effectively bridges the synthetic-to-real gap in fracture surface understanding for everyday objects.
Real Fracture on Fractura. We further evaluate performance across three fracture types. To isolate the impact of unseen objects, $\textsc{Garf}{}_{\text{LoRA}}$ is fine-tuned separately on synthetic fractures from bones, eggshells, and lithics in FracturaFine-tuning is not performed for ceramics, as its object categories closely resemble those in the Everyday subset of Breaking Bad.. Figure 5 compares reassembly performance across three fracture types in the Fractura real fracture subset. Garf outperforms competing methods on ceramics even with three missing fragments. $\textsc{Garf}{}_{\text{LoRA}}$ further improves performance by mitigating the effect of unseen objects. $\textsc{Garf}{}_{\text{LoRA}}$ generalize well to random breakage (limb bones, ceramics) and incomplete ossification (vertebrae). Unfortunately, although finetune significantly improves the performance on lithics, both Garf and $\textsc{Garf}{}_{\text{LoRA}}$ struggle on lithics, likely due to high ambiguity among flakes and the core, a well-known challenging spatial reasoning problem for anthropologists [laughlin_experimental_2010]. This unresolved challenge in Fractura offer valuable directions for future research.
Table 3: Quantitative Results on Fantastic Breaks Dataset [lamb2023fantastic]. This includes manually collected real-world objects.
Methods RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$ CD $\downarrow$
degree $\times 10^{-2}$ % $\times 10^{-3}$
Tested on the Fantastic Breaks
Jigsaw [lu2024jigsaw] 26.30 6.43 73.64 10.47
PF++ [wang2024puzzlefusion++] 20.68 4.37 83.33 6.68
Garf 10.62 2.10 91.00 2.12
Figure 5: Qualitative Comparisons on the Fractura real fracture subset. Garf generalizes well to random breakage (limb bones and ceramics) and incomplete ossification (vertebrae) but faces challenges with high-ambiguity fractures like flintknapping (lithics). Fine-tuning enhances performance, particularly for thin-shell structures (eggshells) and flintnapping (lithics).
4.4 How Do Missing or Extraneous Parts Affect Performance?
Archaeological materials are often incomplete or mixed with similar but extraneous fragments, posing significant challenges for assembly. To quantitatively assess model robustness under these conditions at scale, we extend Breaking Bad’s Everyday subset in two ways: (i) Missing parts subset removes 20% of fragments in descending order of volume, preserving the largest anchor fragments and maintaining the object’s connectivity graph; (ii) Extraneous parts subset adds fragments from other objects in the same category, selecting pieces smaller than the largest anchor fragment but larger than 5% of the object’s total volume to ensure they are not trivially small. For a fair comparison, we evaluate objects with 5 to 17 fragments. Additionally, we provide visual demonstrations of assemblies with naturally missing or extraneous parts in Fractura.
Results and Analysis. Table 4 shows that Garf demonstrates strong resilience, with minimal performance degradation over competing methods. With 20% extraneous fragments, Garf maintains a high PA of 79.21%, only 5.0% lower than with complete sets, whereas PF++ drops by 24.32%. Similarly, with 20% missing fragments, Garf achieves 78.87% PA, far surpassing Jigsaw (28.85%) and PF++ (47.22%). Figure 6 illustrates how missing or extraneous fragments affect reassembly performance. While all methods degrade under these challenging conditions, Garf demonstrates superior robustness, maintaining coherent structures despite missing or extraneous fragments. In contrast, Jigsaw and PF++ exhibit severe misalignments and fragment mismatches. This robustness suggests that Garf can partially handle missing or extraneous fragments, benefiting from our model design.
Table 4: Quantitative Results on Missing / Extraneous Parts.
Methods Input RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$ CD $\downarrow$
Jigsaw [lu2024jigsaw] Complete 71.41 15.83 28.34 21.94
20% Miss. 70.45 15.54 28.85 21.58
20% Extra. 74.55 19.02 24.03 26.15
PF++ [wang2024puzzlefusion++] Complete 59.26 12.00 49.38 5.52
20% Miss. 61.14 12.26 44.71 7.57
20% Extra. 61.25 13.91 40.59 8.77
Garf Complete 19.55 3.83 83.39 0.62
20% Miss. 22.23 4.75 78.87 1.40
20% Extra. 22.70 4.62 79.21 1.29
Figure 6: Qualitative Comparisons on the Missing or Extraneous Impact. Garf demonstrates superior robustness, maintaining coherent structures despite missing or extraneous fragments.
4.5 Ablation Study
Designs of Model. We incrementally add our key design choices to the PF++ [wang2024puzzlefusion++] baseline, reporting RMSE (R/T) and PA as shown in Table 5. 1 Fracture-aware Pretraining Strategy. We replace the VAE-based pretraining used in PF++ with our fracture-aware pretraining strategy. Our strategy reduces RMSE(R) by 53.3% and RMSE(T) by 49.0%. 2 More attention on large fragments. We observe that large fragments are easier to assemble and yield more stable gradients. To leverage this, we apply surface-area-based weighted sampling and modify the attention mechanism to emphasize large fragments. This leads to substantial improvements. 3 Two-Session Flow Matching on SE(3). We design a two-session flow matching scheme: the first session performs one-step coarse pose initialization, while the second refines it via multi-step flow. This design further boosts performance during inference. More experimental results and analysis are in the supplementary materials.
Table 5: Ablation Study on Our Designs of Model.
Setups Backbone Denoiser RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$
I: PuzzleFusion++ PointNet++ Diffusion 32.91 5.26 78.95
I + 1 PointNet++ Diffusion 15.36 2.68 89.40
I + 1 PointNet++ FM 15.55 2.69 89.43
I + 1 2 PointNet++ Diffusion 12.58 2.16 91.24
I + 1 2 PointNet++ FM 8.70 1.67 93.60
I + 1 2 3 PointNet++ FM 7.40 1.47 94.43
I + 1 2 3 PTv3 FM 6.68 1.34 94.77
(GARF-mini)
Figure 7: Investigation on the Number of Fragments. Garf-mini significantly outperforms PF++, with the two-session FM further boosting results for >10>10 fragments. Notably, Garf-mini also generalizes to fractures containing well over 20 fragments.
Number of Fragments. We analyze performance across varying fragment counts, as shown in Figure 7. Garf-mini consistently surpasses PF++ across all fragment counts. Our two-session FM further enhances performance on unseen objects containing more than 10 fragments. Interestingly, although Garf has only been trained on data with 2-20 fragments in Breaking Bad, it can still generalize to fractures with more than 20 pieces as shown in Figure 7.
5 Impact and Limitations
Scientific Impacts. How objects—whether they are bones, ceramic pots, or stone tools—were reassembled and what processes influenced their reconstruction is one of the basic questions common to different research communities, including paleontology and paleoanthropology, archaeology, and forensic science. To explore this, we make the first attempt to collaborate with archaeologists, paleoanthropologists, and ornithologists to build a generalizable model for real-world fracture reassembly. While Garf achieves significant improvements, challenges remain, particularly in handling the complexities presented by Fractura, which creates new opportunities for the vision and learning communities, encouraging advancements in 3D puzzle solving.
Limitations and Future Directions. Garf encounters challenges in handling fracture ambiguity, especially when fragments have subtle geometric differences or inherently ambiguous fracture surfaces. For instance, it struggles with lithic refitting in Fractura due to high ambiguity among flakes and the core, as well as fresco reconstruction in RePAIR, where erosion affects fracture surfaces [tsesmelis2025re]. Therefore, our future work will focus on: (i) Multimodal fracture reassembly, integrating geometric and texture information; (ii) Test-time policy optimization using expert feedback; (iii) Expanding the size and diversity of Fractura.
6 Related Work
Fracture Assembly. Early methods relied on explicit geometric matching with handcrafted features [luo2013co, andreadis2014facet, xu2015robust, holland2022digital], often struggling with ambiguous or incomplete geometries. The advent of the large-scale synthetic dataset Breaking Bad [sellan2022breaking] has enabled learning-based approaches to acquire robust geometric representations and assembly strategies [lu2024survey]. Jigsaw [lu2024jigsaw] jointly learns hierarchical features from global and local geometries for fracture matching and pose estimation, while SE(3)-equiv [wu2023leveraging] extracts fragment features for pose estimation. DiffAssemble [scarpellini2024diffassemble] improves performance using a diffusion model. PuzzleFusion++ [wang2024puzzlefusion++] mimics how humans solve spatial puzzles by integrating diffusion-based pose estimation with a VAE-based fragment representation and transformer-based alignment verification. However, while PuzzleFusion++ achieves SOTA results on the everyday subset, its performance degrades significantly on the Artifact subset [wang2024puzzlefusion++]. More critically, it remains unclear whether models trained on synthetic data can generalize to real-world fractures with more complex breakage patterns. To fill this gap, we identify major real-world fracture challenges and curate Fractura, a dataset capturing key complexities. To address these challenges, we propose Garf, a generalizable 3D reassembly framework for real-world fractures.
Flow Matching. Recent advances have explored flow matching (FM) across image generation [esser2024scaling], protein backbone generation [yim2023fast, huguet2024sequence, geffnerproteina], and general robot control [black2024pi0]. AssembleFlow [guo2025assembleflow] attempts to leverage FM for molecular assembly, but introduces numerical errors by approximating quaternion updates through direct addition over small time intervals during inference. While diffusion models have been widely applied to fracture assembly [scarpellini2024diffassemble, wang2024puzzlefusion++], FM provides a more natural formulation by learning geodesic flows in SE(3). We propose the first FM-based fracture assembly framework, incorporating the multi-anchor training strategy and two-session flow matching at inference time.
7 Conclusion
In collaboration with archaeologists, paleoanthropologists, and ornithologists, we present Fractura, a diverse and challenging fracture assembly dataset, and Garf, a generalizable 3D reassembly framework designed for real-world fractures. Fractura serves as a challenging benchmark to evaluate how object shapes, fracture types, and the presence of missing or extraneous parts affect reassembly performance. Facing these challenges, Garf offers vital guidance on training on synthetic data to advance real-world 3D puzzle solving. Comprehensive evaluations demonstrate its strong generalization to unseen object shapes and diverse fracture types. Despite its superior performance, Garf still struggles with geometric ambiguity, particularly when dealing with highly similar fragments and eroded fracture surfaces. We anticipate that Fractura will drive further advancements in 3D reassembly, pushing the boundaries of spatial reasoning to answer unknown scientific questions.
Acknowledgement. We gratefully acknowledge the Physical Anthropology Unit, Universidad Complutense de Madrid for access to curated human skeletons, and Dr. Scott A. Williams (NYU Anthropology Department) for the processed data samples. This work was supported in part through NSF grants 2152565, 2238968, 2322242, and 2426993, and the NYU IT High Performance Computing resources, services, and staff expertise.
Appendix
This document supplements the main paper as follows:
  1. 1.
    Dataset details (Section A).
  2. 2.
    More details about the training recipe and reproducibility (section B).
  3. 3.
    More visualizations and detailed tables (section C).
A Additional Dataset Details
A.1 Fracture Simulation
(i) Bone. For elongated structures like limbs and ribs, we used Blender’s skinning and subdivision surface techniques to create realistic cylindrical hollows, replicating bone morphology. We then applied the physics-based fracture method from Breaking Bad [sellan2022breaking] to generate 2–20 fragments. The same approach was used for os coxae and vertebrae, forming the simulated subset of the bone category. (ii) Eggshell. Since scanned eggshells produce watertight solid ellipsoids, we removed 98% of the concentric volume to simulate thin shells. We then applied the same physics-based fracture method to generate realistic breakage patterns. (iii) Ceramics. Given that ceramic objects (e.g., bowls, pots, vases) closely resemble those in Breaking Bad’s everyday category, we focused on scanning real fragments and did not include a simulated subset. (iv) Lithics. As an initial feasibility test, two generalized core morphologies were repeatedly virtually knapped with some randomized variation following methods described for the dataset in [orellana2021proof] to produce core and flake combinations with varying geometries.
A.2 Fractura Statistics
Table I presents detailed statistics for each category in Fractura. We continue to expand both the dataset’s scale and diversity, aiming to establish a comprehensive cyberinfrastructure for the vision-for-science community.
Table I: Dataset Statistics of the Fractura Dataset.
Category Fracture Type # Assemblies # Pieces
Bone Real 17 37
Synthetic 7056 39943
Eggshell Real 3 12
Synthetic 2268 12600
Ceramics Real 9 51
Synthetic N/A N/A
Lithics Real 12 192
Synthetic 403 807
Total Real 41 292
Synthetic 9727 53350
B Additional Implementation Details
B.1 Data Preprocessing
We preprocess the BreakingBad dataset [sellan2022breaking] to calculate the segmentation ground truth directly from meshes to reduce the computation overhead during training as described in Sec. 3.1, and there’s no need for any hyperparameters. Unlike baseline methods (Global, LSTM, and DGL) provided by the dataset and PF++ [wang2024puzzlefusion++], which samples $M=1000$ points from the mesh per fragment, we used the same setting as in Jigsaw [lu2024jigsaw] to sample $M=5000$ points per object, making all fragments have the same point density. With this sampling setting, we did not encounter any gradient explosion issues during training, as reported in FragmentDiff [xu2024fragmentdiff], which occur when sampling too many points for tiny pieces. Meanwhile, we employ the Poisson disk sampling method to ensure that the points are more uniformly distributed on the surface of the fragment. During training, standard data augmentation techniques are applied, including recentering, scaling, and random rotation.
B.2 Training Recipe
We modified a smaller version of Point Transformer V3 [wu2024point] as our backbone for the segmentation pretraining, as shown in Table II, which we found to be sufficient and more memory efficient. Since Garf uses a much larger training dataset, we reduce the training epochs to 150, other than the 400 epochs used in Garf-mini. Both pretrainings reach over 99.5% accuracy on the validation set. Samples of segmentation results are shown in Fig. I.
Figure I: Segmentation results on a real-world object (left), Breaking Bad [sellan2022breaking] (center) and Fantastic Breaks [lamb2023fantastic] (right).
For FM training, we provide the hyperparameters in Table II for reproducibility. The settings are identical for both Garf and Garf-mini, as their only difference lies in the pretraining stage.
Table II: Training Configurations.
Config Value
Backbone Encoder Depth [2, 2, 6, 2]
Encoder # Heads [2, 4, 8, 16]
Encoder Patch Size [1024, 1024, 1024, 1024]
Encoder Channels [32, 64, 128, 256]
Decoder Depth [2, 2, 2]
Decoder # Heads [4, 8, 16]
Decoder Patch Size [1024, 1024, 1024]
Decoder Channels [256, 128, 64]
Pretraining Global Batch Size 256
Epochs 400 / 150
Learning Rate 1e-4
Scheduler CosineAnnealingWarmRestarts
Scheduler $T_{0}$ 100 / 50
# Trainable Params 12.7M
Training Global Batch Size 128
Epochs 1500
Learning Rate 2e-4
Scheduler MultiStepLR
Scheduler Milestones [900, 1200]
Scheduler $\gamma$ 0.5
# Trainable Params 43.5M
B.3 Preliminaries on Riemannian Flow Matching
Instead of simulating discrete noise addition steps, flow matching (FM) learns a probability density path $p_{t}$, which progressively transforms a noise distribution $p_{t=0}$ to the data distribution $p_{t=1}$, with a time variable $t\in[0,1]$. As a simulation-free method aiming to learn continuous normalizing flow (CNF), FM models a probability density path $p_{t}$, which progressively transforms a noise distribution $p_{t=0}$ to the data distribution $p_{t=1}$, with a time variable $t\in[0,1]$. Inspired by learning assembly by breaking, the rigid motion of the fragments corresponds to the geodesic on the Lie group $\mathrm{SE}(3)$, which is a differentiable Riemannian manifold. Inspired by previous works [bose2023se, yim2023fast, geffnerproteina], FM can be extended to $\mathrm{SE}(3)$ manifold to learn the rigid assembly process.
On a manifold $\mathcal{M}$, the flow $\psi_{t}:\mathcal{M}\rightarrow\mathcal{M}$ is defined as the solution of an ordinary differential equation (ODE):
$\frac{\mathrm{d}}{\mathrm{d}t}\psi_{t}(\bm{x})=\bm{v}_{t}(\psi_{t}(\bm{x})),\quad\psi_{0}(\bm{x})=\bm{x},$ (8)
where $\bm{v}_{t}(\bm{x})\in\mathcal{T}_{\bm{x}}\mathcal{M}$ is the time-dependent vector field, and $\mathcal{T}_{\bm{x}}\mathcal{M}$ is the tangent space of the manifold at $\bm{x}\in\mathcal{M}$. In the context of $\mathrm{SE}(3)$, the tangent space is the Lie algebra $\mathfrak{se}(3)$, which is a six-dimensional vector space, presenting the velocity of the rigid motion of the fragments. Given the conditional vector filed $\bm{u}_{t}(\bm{x}\mid\bm{x}_{1})\in\mathcal{T}_{\bm{x}}\mathcal{M}$, which generates the conditional probability path $p_{t}(\bm{x}\mid\bm{x}_{1})$, the Riemannian flow matching objective can be defined as:
$\mathcal{L}_{\text{CFM}}:=\mathbb{E}_{t,p_{1}(\bm{x}_{1}),p_{t}(\bm{x}\mid\bm{x}_{1})}\left[\|{\bm{v}_{t}(\bm{x},t)-\bm{u}_{t}(\bm{x}\mid\bm{x}_{1})}\|^{2}_{G}\right],$ (9)
where $\|\cdot\|^{2}_{G}$ is the norm induced by the Riemannian metric $G$. Then the learned vector field $\bm{v}_{t}$ can be used to generate samples on the manifold at inference, which is $\mathrm{SE}(3)$ poses of the fragments. The rigid motion of fragments corresponds to the geodesic on the Lie group $\mathrm{SE}(3)$, a differentiable Riemannian manifold.
B.4 More attention on large fragments
Garf provides tailored designs to place more attention on large fragments. We observed that: (i) large fragments are easier to assemble; (ii) tiny fragments sometimes lead to unstable training. Driven by these insights, we apply weighted sampling based on the surface area of fragments and modify the self-attention module to allow more attention on large fragments.
Figure II: Self-attention comparison between Garf (left) and PuzzleFusion++ [wang2024puzzlefusion++] (right).
Table III: Results on Vanilla Breaking Bad [sellan2022breaking] Dataset.
Methods RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$ CD $\downarrow$
degree $\times 10^{-2}$ % $\times 10^{-3}$
Tested on the Everyday Subset
Global [li2020learning] 80.70 15.10 24.60 14.60
LSTM [wu2020pq] 84.20 16.20 22.70 15.80
DGL [zhan2020generative] 79.40 15.00 31.00 14.30
SE(3)-Equiv [wu2023leveraging] 79.30 16.90 8.41 28.50
DiffAssemble [scarpellini2024diffassemble] 73.30 14.80 27.50 -
PHFormer [cui2024phformer] 26.10 9.30 50.70 9.60
Jigsaw [lu2024jigsaw] 42.30 10.70 57.30 13.30
PF++ [wang2024puzzlefusion++] 38.10 8.04 70.60 6.03
Garf-mini 10.41 1.91 92.77 0.45
Tested on the Artifact Subset
Jigsaw 52.40 22.20 45.60 14.30
PF++ 52.10 13.90 49.60 14.50
Garf-mini 11.91 2.74 89.42 1.05
Table IV: Ablation Study on Our Designs of FM.
$\mathrm{SE}(3)$ Multi-Anchor One-Step RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$
10.24 1.95 89.08
8.02 1.63 93.78
7.63 1.60 94.02
6.68 1.34 94.77
Table V: Ablation Study on Sample Steps.
Steps RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$ CD $\downarrow$ Speed (ms)
1 12.52 3.18 86.88 2.14 38.26
One-Step + 1 9.79 2.46 91.31 1.42 45.76
5 8.25 1.92 93.70 0.53 57.32
One-Step + 5 7.15 1.66 94.43 0.46 76.23
20 7.63 1.60 94.02 0.35 185.05
One-step + 20 6.68 1.34 94.77 0.25 190.77
50 7.50 1.54 94.01 0.32 408.40
Table VI: Ablation Study on the Different Anchor Initialization.
Settings RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$ CD $\downarrow$
Largest Anchor 6.10 1.22 95.33 0.22
Random Anchor 6.09 1.30 95.20 0.29
Anchor-Free 9.09 2.13 93.23 0.91
Table VII: Comparison Between Diffusion and Our FM Models.
Dataset Methods RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$
Everyday Diffusion 7.45 1.47 94.30
$\mathrm{SE}(3)$ Diffusion N/A N/A N/A
Diffusion w/ One-Step 7.51 1.47 94.27
Vanilla FM 10.24 1.95 89.08
Garf-mini 6.68 1.34 94.77
Fractura Diffusion $32.38$ $7.90$ $71.73$
Garf-mini 27.88 6.79 76.25
C Additional Results and Analyses
C.1 Ablation on Design Choices in Flow Matching
We conduct an ablation study to evaluate the impact of design choices in our FM module. As shown in Table IV, vanilla FM, trained with spherical linear interpolation (slerp) to approximate valid rotations in the forward process [guo2025assembleflow], achieves 89.08 PA, already surpassing previous methods [wang2024puzzlefusion++, lu2024jigsaw]. Incorporating the $\mathrm{SE}(3)$ representation further improves performance by pre-modeling the manifold distribution and better capturing distribution shifts during assembly. Multi-anchor training strategy further enhances results, while one-step pre-assembly significantly boosts performance by providing a more reasonable initial pose distribution, leading to the best overall outcomes.
C.2 Ablation on Sample Steps
Table V shows the effect of varying sampling steps in our framework. Surprisingly, even with just 5 steps, FM achieves 93.70% PA, highlighting its effectiveness in modeling global probabilistic paths. Additionally, our first-session initialization provides a more reasonable initial pose, further improving assembly quality while adding minimal computational overhead.
C.3 Ablation on Anchor Fragment
Similar to PF++ [wang2024puzzlefusion++], we use the largest fragment as the anchor fragment at inference. We compare the performance of using the largest fragment, a randomly selected fragment, and no anchor fragment. As shown in Table VI, using a random fragment as the anchor fragment has almost no negative effect on the model. Only anchor-free initialization leads to a slight performance drop.
C.4 Comparison with Diffusion Models
Table VII compares our FM module with diffusion models. While diffusion, when paired with fracture-aware pretraining, achieves competitive performance, directly applying vanilla FM yields lower results (89.08 PA), emphasizing the importance of our subsequent design choices. A key limitation of diffusion models is their handling of $\mathrm{SO}(3)$ rotation, which cannot be naturally incorporated into the reverse process. Existing methods, such as score prediction [yim2023se], aim to maintain rotation validity but fall outside our current scope. Additionally, diffusion models rely on multi-step denoising without explicitly modeling the global probabilistic path, rendering one-step pre-assembly ineffective. Furthermore, on Fractura, diffusion models exhibit weaker generalization to unseen objects compared to Garf-mini.
C.5 Quantitative Results on Vanilla Breaking Bad
Given that all our previous experiments were conducted on the volume-constrained version of the Breaking Bad dataset [sellan2022breaking], we here provide additional quantitative results on the non-volume-constrained version to align with the settings of previous methods. The results, shown in Table III, demonstrate that our Garf-mini model still significantly outperforms the previous state-of-the-art method, PF++ [wang2024puzzlefusion++], by a large margin. This performance is consistent across both the everyday and artifact subsets, showcasing the model’s robust generalization ability.
We also present the results of FragmentDiff [xu2024fragmentdiff] on their custom Breaking Bad dataset in Table VIII. FragmentDiff claims to remove tiny pieces, but it is unclear whether this applies only to their training setting or also to evaluation. Unfortunately, since they did not open source their code or provide their preprocessed data, we are unable to directly compare all other methods with FragmentDiff. Additionally, they did not adhere to the common settings used by other methods, which limit the number of pieces from 2 to 20, making direct comparisons on their provided metrics impossible. However, its significant performance drop from the Everyday subset to the Artifact subset suggests that Garf surpasses FragmentDiff in generalization capability.
C.6 Quantitative Results of Finetuning on the Fractura Synthetic Dataset
After finetuning Garf on the Fractura synthetic dataset, we report the per-category performance on the bone and eggshell categories, as shown in Table IX. The results demonstrate that finetuning the FM model in Garf significantly improves performance on these two unseen categories, showing the effectiveness of our finetuning techniques and the generalizability of our pretraining strategy.
C.7 Additional Qualitative Comparison on the Fractura and Breaking Bad Dataset
Figures IIIIV and V demonstrate more qualitative comparison on the Fractura and Breaking Bad Dataset, where our Garf shows superior performance than the other previous SOTA methods.
Table VIII: FragmentDiff [xu2024fragmentdiff] Results on Their Custom Breaking Bad Dataset.
Methods Subset RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$
degree $\times 10^{-2}$ %
FragmentDiff [xu2024fragmentdiff] Everyday 13.68 7.41 90.20
Artifact 18.18 8.12 82.30
Table IX: Quantitative Per-category Results on the Fractura (Synthetic Fracture).
Category Method RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$ CD $\downarrow$
degree $\times 10^{-2}$ % $\times 10^{-3}$
Bone Jigsaw 66.44 20.54 27.24 91.70
PF++ 66.28 20.50 29.81 47.78
Garf 17.70 3.80 85.18 5.11
$\textsc{Garf}{}_{\text{LoRA}}$ 8.79 1.10 98.19 0.34
Eggshell Jigsaw 44.44 12.88 49.03 10.49
PF++ 54.81 13.81 61.36 1.50
Garf 22.48 6.16 83.41 0.67
$\textsc{Garf}{}_{\text{LoRA}}$ 7.10 1.95 95.68 0.26
Figure III: Qualitative Results on the Fractura Synthetic Dataset.
Figure IV: Qualitative Results on the Fractura Synthetic Dataset.
Figure V: Qualitative Results on the Breaking Bad Dataset Artifact Subset.